OpenStudy (anonymous):

a package dropped from a helicopter falls 4.7m during the first second, 13.6m during the second, 22.5m during the third second and so on. how high is the helicopter if the package takes 10 seconds to reach the ground?

OpenStudy (anonymous):

hint: arithmetic series

OpenStudy (anonymous):

yes I found it to be arithmetic

OpenStudy (anonymous):

I am stuck

OpenStudy (anonymous):

do you know how to sum arithmetic series?

OpenStudy (anonymous):

Sn=n/2[2a1+(n-1)d]

OpenStudy (anonymous):

so can you identify \(a_1\), \(n\) and \(d\)?

OpenStudy (anonymous):

a1=4.7 and n=10 d=8.9

OpenStudy (anonymous):

yep, so now you can calculate answer yourself?

OpenStudy (anonymous):

and to clarify this is what you write for sum formula? \[S_n=\frac{2a_1+(n-1)d}{2}n\] right?

OpenStudy (anonymous):

I will do it then I will ask you to check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first step: an=a1+(n-1)d a10=4.7+(10-1)8.9=84.8

OpenStudy (anonymous):

shouldnt it be 2a1?

OpenStudy (anonymous):

did I use wrong formule

OpenStudy (anonymous):

ah you calculating \(a_n\)? why?

OpenStudy (anonymous):

second step: Sn=n/2(a1+an) Sn=10/2(4.7+84.8) Sn=447.5

OpenStudy (anonymous):

yeh you can do that way too and your answer is correct

OpenStudy (anonymous):

how can i do it without calculating an?

OpenStudy (anonymous):

you have written yourself: Ocuane Sn=n/2[2a1+(n-1)d] 21 minutes ago so you can just plug in numbers but it's actually same but you could do in 1 step

OpenStudy (anonymous):

just hard to judge which formule to use

OpenStudy (anonymous):

its same formula

OpenStudy (anonymous):

\[S_n=\frac{a_1+a_n}{2}n=\frac{a_1+a_1+(n-1)d}{2}n=\frac{2a_1+(n-1)d}{2}n\]

OpenStudy (anonymous):

if you know \(a_n\) you can use first one but if you don't you can use second one without even evaluating \(a_n\)

OpenStudy (anonymous):

so first I have to identify what i have to decide which formule to use am i right?

OpenStudy (anonymous):

yeah but you can always use one formula it just save a bit of time

OpenStudy (anonymous):

I tryed to plug in without calculating an and did not get the same result

OpenStudy (anonymous):

try it again you should get same result

OpenStudy (anonymous):

I tryed and got 127.1

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

Sn=n/2[2a1+(n-1)d] =10/2[2*4.7+(10-1)8.9] =127.1

OpenStudy (anonymous):

\[\frac{2\cdot4.7+(10-1)8.9}{2}10\] did you write like ths?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so this is correct formula

OpenStudy (anonymous):

do you understand now?

OpenStudy (anonymous):

still doing it

OpenStudy (anonymous):

yes got it had done calculations wrong thanks so much